### v1

# : A Quantum Multiple Kernel Learning Approach for Drug Discovery

Abstract
Quantitative Structure-Activity Relationship () modeling is a cornerstone of computational drug discovery. This research demonstrates the successful application of a Quantum Multiple Kernel Learning () framework to enhance classification, showing a notable performance improvement over classical methods. We apply this methodology to a dataset for identifying kinase inhibitors. The workflow involves converting representations into numerical molecular descriptors, reducing dimensionality via Principal Component Analysis (), and employing a Support Vector Machine () trained on an optimized combination of multiple quantum and classical kernels. By benchmarking the - against a classical Gradient Boosting model, we show that the quantum-enhanced approach achieves a superior score, highlighting its potential to provide a quantum advantage in challenging cheminformatics classification tasks.
Index Terms:
, classification, drug discovery, quantum machine learning, multiple kernel learning, support vector machines.I Introduction
II QSAR Theory and Contemporary Relevance
II-A Foundations of Quantitative Structure-Activity Relationships
Quantitative Structure-Activity Relationship () modeling represents a cornerstone (one of them) methodology in computational chemistry and drug discovery, embodying the fundamental principle that molecular structure can determine biological activity. This paradigm, first conceptualized by Hansch and Fujita in the 1960s, has evolved into a sophisticated computational framework that bridges chemical structure and pharmacological function through mathematical modeling [1].
The theoretical foundation of rests on the premise that similar molecular structures exhibit similar biological activities a principle that enables the systematic prediction of pharmacological, toxicological, and physicochemical properties from structural information alone. In essence, transforms the complex relationship between molecular architecture and biological response into a tractable computational problem through the systematic encoding of chemical structures into numerical descriptors [2].
Structurally, models function as statistical or machine learning frameworks that utilize molecular descriptors as predictor variables for properties of interest, whether physicochemical (solubility, lipophilicity), biological (enzyme inhibition, receptor binding), or toxicological (mutagenicity, carcinogenicity). The general workflow encompasses molecular representation via strings, descriptor calculation, feature engineering, model development, and validationβa pipeline that has become increasingly sophisticated with advances in cheminformatics and machine learning.
II-B Contemporary Relevance and Industrial Impact
The relevance of methodologies in modern drug discovery has intensified dramatically due to several converging factors that define the current pharmaceutical landscape. The exponential growth in chemical databases, with repositories like ChEMBL and PubChem containing millions of bioactive compounds, necessitates computational approaches capable of rapid property prediction and compound prioritization.
Three fundamental advantages position as an indispensable tool in contemporary drug development:
(i) Predictive Efficiency: Once developed and validated, models enable instantaneous property prediction from molecular structure alone, eliminating the need for costly and time-intensive experimental screening. This capability is particularly valuable in early-stage drug discovery, where rapid assessment of large compound libraries is essential for lead identification and optimization.
(ii) Scalability and Automation: Modern frameworks support fully automated evaluation of massive chemical spaces, facilitating the assessment of millions of virtual compounds through high-throughput computational screening. This scalability aligns with current trends toward big data approaches in pharmaceutical research, where comprehensive exploration of chemical space is increasingly feasible.
(iii) Regulatory Acceptance: models have achieved widespread acceptance by regulatory agencies, including the , , and , particularly for toxicity assessment and environmental risk evaluation. This regulatory validation ensures that -based predictions can support formal safety assessments and regulatory submissions, providing a pathway from computational prediction to clinical application.
II-C Machine Learning Enhancement and Quantum Frontiers
The integration of advanced machine learning algorithms into workflows has revolutionized the fieldβs predictive capabilities, enabling the capture of complex, non-linear structure-activity relationships that were previously intractable. Deep learning architectures, ensemble methods, and sophisticated feature selection algorithms have expanded βs applicability to increasingly challenging molecular targets and property predictions.
However, as chemical datasets grow in complexity and dimensionality, traditional computational approaches encounter fundamental limitations in feature space exploration and pattern recognition. This challenge motivates the exploration of quantum machine learning approaches, which leverage the exponentially large Hilbert spaces of quantum systems[3], to potentially identify molecular patterns that remain hidden to classical algorithms.
The emergence of Quantum Machine Learning () in cheminformatics represents a paradigm shift toward harnessing quantum computational principles for enhanced molecular property prediction.
By mapping classical molecular data into quantum states through specialized feature maps, quantum algorithms can potentially access richer representational spaces and capture subtle structure-activity relationships that classical methods might overlook [4].
Quantitative Structure-Activity Relationship () models aim to establish a quantitative link between the chemical structure of a molecule and its biological activity [5]. These mathematical models provide an in silico methodology to predict the properties of novel compounds, thereby prioritizing experimental efforts and reducing costs. Applications range from predicting pharmacokinetic processesβabsorption, distribution, metabolism, and excretion ()βto assessing toxicological endpoints, and even modelling new properties
The evolution of has seen a shift from simple linear regression to more sophisticated machine learning algorithms. Methods like Random Forests [6] and Bayesian Neural Networks [7] have shown superior performance by capturing complex, non-linear relationships within the data. However, as molecular data grows in complexity, in the sense its important to describe our compound in more ways than just than a few feature our dataset that captures all this relations growths also in complexity. There is an ongoing need for methods that can effectively navigate high-dimensional feature spaces.
Quantum Machine Learning () offers a promising new direction. By mapping classical data into the exponentially large Hilbert spaces of quantum systems, algorithms, particularly Quantum Support Vector Machines () [8], have the potential to identify patterns that are intractable for classical models [9]. This work focuses on a specific enhancement of known as Quantum Multiple Kernel Learning () [10], where we combine the strengths of various quantum and classical kernels to build a more robust and accurate classifier for a real-world drug discovery problem.
III Methodology
Our pipeline consists of three main stages: (1) data pre-processing and feature engineering, (2) classification using a - model, and (3) benchmarking against a classical Gradient Boosting model.
III-A Dataset and Preprocessing
The study utilizes a public dataset of (Dual-specificity tyrosine-phosphorylation-regulated kinase 1A) inhibitors [11].
The dataset was originally shared by a second research group, which made the model classically, so it is our starting point for performing a benchmark against the same dataset.
A comprehensive dataset of 1291 molecules exhibiting inhibitory activity against human kinase was systematically compiled from the database (). represents a critical therapeutic target for neurodegenerative disorders, particularly Alzheimerβs disease, due to its role in tau phosphorylation and neuronal dysfunction [12].
The kinaseβs involvement in Down syndrome-associated cognitive impairment and its potential as a druggable target for neuroprotective interventions makes it an ideal case study for quantum-enhanced modeling approaches.
The raw dataset underwent rigorous quality control procedures to ensure data integrity and experimental consistency. The curation pipeline comprised three critical stages: (i) identification and removal of duplicate molecular entries, (ii) validation of chemical structures to eliminate invalid representations, and (iii) application of bioassay selection criteria. Specifically, only compounds from experiments annotated with BioAssay Ontology () label βsingle protein assayβ were retained to ensure direct target-compound interaction measurements, excluding complex cellular or multi-target assays that could introduce confounding factors.
Following systematic curation, the final data set comprised 354 unique molecules with experimentally determined values that ranged from 0.5 to 100 M. For binary classification, compounds were categorized using a standard medicinal chemistry threshold established for kinase inhibitors: molecules with (corresponding to ) were classified as βactiveβ inhibitors, while those below this threshold were designated as βinactiveβ. [13].
III-A1 Molecular Descriptors
The feature set comprised representative molecular descriptors from established cheminformatics libraries ().
Shape and Geometric Descriptors
PMI1 (Principal Moment of Inertia, 1st order): The first principal moment of inertia calculated from the 3D molecular coordinates, representing the distribution of atomic masses around the principal axis. PMI1 values typically range from 10 to 1000 amuΓ 2 for drug-like molecules, with higher values indicating more elongated structures.
NPR1 (Normalized Principal Ratio 1): The ratio PMI1/PMI3, normalized to eliminate molecular size effects. NPR1 values range from 0.5 (spherical) to 1.0 (linear), providing size-independent shape characterization essential for scaffold-hopping applications
Electronic and Topological Descriptors
MolLogP (Molecular LogP): Octanol-water partition coefficient calculated using the Wildman-Crippen method implemented in . Values typically range from -3 to +8 for drug-like compounds, with optimal range 1-3 for oral bioavailability
TPSA (Topological Polar Surface Area): Sum of surfaces of polar atoms (oxygen and nitrogen) in a molecule, calculated using group contributions. TPSA values Β‘140 Γ Β² correlate with good blood-brain barrier permeability
NumRotatableBonds: Count of rotatable bonds in the molecule, excluding amide bonds. Drug-like molecules typically have 10 rotatable bonds according to Lipinskiβs rule
Connectivity and Graph-Based Descriptors
BertzCT (Bertz Complexity Index): A graph-theoretical descriptor that quantifies molecular complexity by considering both the number of atoms and their connectivity patterns. Typical values range from 0 for simple molecules (e.g., methane) to over 1000 for complex natural products [14]. Its considers both the number of atoms and their connectivity patterns within a molecular structure. This index helps in understanding and predicting the complexity of molecules, particularly in the context of synthetic planning and analysis of synthetic strategies. The Bertz Complexity Index utilizes graph theory to represent molecules. In this representation, atoms are nodes (vertices) and bonds are edges within the graph
Chi0n, Chi1n (Kier-Hall Connectivity Indices): Topological indices describing molecular branching and connectivity. Chi0n reflects molecular size, while Chi1n captures first-order connectivity patterns
Pharmacophore and Functional Group Descriptors
NumHBD (Hydrogen Bond Donors): Count of hydrogen bond donor groups (OH, NH). Drug-like molecules typically have 5 donors
NumHBA (Hydrogen Bond Acceptors): Count of hydrogen bond acceptor atoms (N, O). Optimal range 10 for oral drugs
FractionCsp3: Fraction of carbons that are sp3 hybridized. Higher Fsp3 values () correlate with increased three-dimensionality and improved developability
NumAromaticRings: Count of aromatic ring systems in the molecule. Excessive aromatic content ( rings) often correlates with poor solubility and toxicity issues.
Validation and Quality Control
All descriptors were calculated using 2023.03.1 with default parameters and validated against reference compounds from the database. Descriptor values were checked for reasonable ranges based on drug-like chemical space: molecular weight (150-500 Da), LogP (-2 to 5), TPSA (20-140 Γ 2), and rotatable bonds (0-10). This validation ensures that the quantum feature mapping operates on chemically meaningful and computationally stable molecular representations.
This high-dimensional descriptor matrix provides a comprehensive mathematical representation of molecular structure, capturing diverse aspects of chemical information essential for robust QSAR model development
The dataset is pre-divided into a training set and a test set, with each compound labeled with a binary target, , indicating its inhibitory activity.
A multi-step preprocessing pipeline was implemented to prepare the data for our models:
1.
Numerical Feature Extraction: The primary molecular structure is given as a (Simplified Molecular Input Line Entry System) string. As machine learning models require numerical input, we used the library to convert each string into a high-dimensional vector of approximately 200 physicochemical descriptors (e.g., molecular weight, LogP, topological polar surface area).
2.
Standardization: The resulting numerical feature set was standardized using Scikit-learnβs StandardScaler, which scales each feature to have a mean of 0 and a standard deviation of 1. The scaler was fitted only on the training data to prevent data leakage.
3.
Dimensionality Reduction: To create a feature set suitable for current quantum hardware and simulators, was applied. We reduced the high-dimensional descriptor space to principal components, which capture the maximum variance in the data. These four components serve as the final input features for all subsequent models.
III-B Quantum Multiple Kernel Learning
Support Vector Machines () are powerful classifiers that work by finding an optimal hyperplane separating data points in a high-dimensional feature space. The βkernel trickβ allows s to operate in this space implicitly, without ever computing the coordinates of the data in that space. A kernel function, , computes the similarity between two data points and .
Our approach extends this concept by using Quantum Multiple Kernel Learning [10].
III-B1 Quantum Kernels
A quantum kernel, , is created by first encoding classical data points into quantum states via a quantum feature map, which is a parameterized quantum circuit [15]. The kernel value is then the similarity measure between these two quantum states, typically the fidelity:
This process leverages quantum properties like superposition and entanglement to potentially map the data into a richer feature space [16]. We used a feature map composed of angle-embedding rotations and entangling gates [17].
III-B2 Kernel combination
Instead of relying on a single kernel, seeks to find the best possible linear combination of a pre-defined set of basis kernels given by
In our work, the set of basis kernels included several different quantum fidelity kernels (generated by varying the feature map parameters) and a classical Radial Basis Function (RBF) kernel. The optimal weights were determined using a average alignment, which maximizes the similarity between the combined kernel and an ideal target kernel derived from the training labels.
III-C Classical Benchmark Model
A classical from Scikit-learn was trained and evaluated on the exact same 4-dimensional, -processed data. This ensures a fair and direct comparison of the classification power of the against a strong, widely-used classical algorithm.
IV Results and Discussion
The performance of both the final model and the classical Gradient Boosting model was evaluated on the unseen test set. The primary metric for comparison is the Area Under the Receiver Operating Characteristic Curve ().
Table I summarizes the final performance scores. The model achieved an of 0.8750, demonstrating a clear and significant performance improvement over the classical Gradient Boosting model, which scored an of 0.8037. This result highlights the effectiveness of the quantum kernel-based approach for this classification task.
| Model | ROC-AUC Score |
| QMKL-SVM | 0.8750 |
| Gradient Boosting | 0.8037 |
Figure 1 provides a visual comparison of the curves for both models. The plot clearly illustrates the superior performance of the , as its curve consistently stays above the Gradient Boosting curve, thereby encompassing a larger area. Both models perform significantly better than a random classifier, but the quantum-enhanced model shows a distinct advantage in its ability to separate the active and inactive compounds.
The superior performance of the model suggests that the combination of multiple quantum kernels and a classical kernel was able to capture more complex, non-linear relationships within the dataβs feature space than the Gradient Boosting algorithm alone. While the computational cost of simulating quantum kernels is currently a limitation, these results strongly support the potential for quantum machine learning to provide a tangible advantage in real-world scientific applications.
V Conclusion
In this work, we successfully developed and benchmarked a complete pipeline for a classification task, demonstrating that a quantum-enhanced model can outperform a strong classical baseline. Our approach, which combines quantum and classical kernels, achieved a higher score than a tuned Gradient Boosting model on the inhibitor dataset.
The workflow, from conversion with to dimensionality reduction with and finally to classification with a quantum-enhanced , represents a viable and powerful strategy for applying to drug discovery. The results indicate that by mapping data into quantum Hilbert spaces, we can unlock new predictive capabilities. As quantum hardware continues to mature, fully leveraging the power of these quantum kernel methods will become increasingly feasible, positioning as a valuable and impactful tool for the future of computational chemistry and pharmacology.
It is important to highlight in this landscape that further investigations are required on how to represent and leverage descriptors in additional dimensions to enhance the robustness and predictive power of models.
β
Acknowledgment
This work was supported by the project funded by Ministerio de Ciencia, InnovaciΓ³n y Universidades, through .